### 07 Apr 2026 v1

# Mixed Hodge Modules and Canonical Perverse Extensions for Multi-Node Conifold Degenerations

Abstract.
We study one-parameter conifold degenerations whose central fiber has finitely many ordinary double points and construct a mixed-Hodge-module refinement of the canonical corrected perverse object associated with the degeneration. We build a rank-one point-supported mixed-Hodge-module block at each node, identify the global singular quotient as
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),$
and assemble these local blocks via Saito’s divisor-case gluing formalism into a global object
$\mathcal{P}^{H}\in MHM(X_{0}).$
We prove that $\mathcal{P}^{H}$ realizes the corrected perverse object, fits into an exact sequence
$0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\to 0,$
and that the same quotient realizes the finite local vanishing sector in the nearby-cycle formalism. We further relate the mixed-Hodge-module extension, its realized perverse extension, and the induced extension on hypercohomology carrying the limiting mixed Hodge structure. This gives a theorem-level Hodge-theoretic refinement of the corrected perverse extension in the finite multi-node ordinary double point setting.
Key words and phrases:
conifold degeneration, perverse sheaves, Picard–Lefschetz theory, spherical twists
2020 Mathematics Subject Classification:
14D06, 32S30, 18G80
Email: arahman@alum.howard.edu
1. Introduction
Let
$\pi:\mathcal{X}\to\Delta$
be a one-parameter degeneration of complex algebraic varieties whose central fiber $X_{0}$ has finitely many ordinary double points
$\Sigma=\{p_{1},\dots,p_{r}\}\subset X_{0}.$
Associated with this degeneration is the canonical corrected perverse sheaf $\mathcal{P}$, obtained from the nearby- and vanishing-cycle formalism. In the single-node case, this object was constructed and characterized in [11] and shown to be the unique minimal Verdier self-dual perverse extension of the shifted constant sheaf across the node in [10]. Additionally, in [10], the construction was placed in the context of limiting mixed Hodge theory and Saito’s nearby-cycle formalism for mixed Hodge modules, thereby isolating the main Hodge-theoretic problem left open by the perverse-sheaf picture: the construction of a fully internal mixed-Hodge-module refinement of the corrected perverse object.
The purpose of the present paper is to solve that problem in the case of finitely many ordinary double points. More precisely, we construct an object
$\mathcal{P}^{H}\in MHM(X_{0})$
whose realization under
$\mathrm{rat}:MHM(X_{0})\to\mathrm{Perv}(X_{0};\mathbb{Q})$
recovers the canonical corrected perverse object $\mathcal{P}$, and whose singular quotient is the finite direct sum of the rank-one point-supported mixed Hodge modules attached to the nodes,
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
We then show that the same mixed-Hodge-module construction governs the vanishing part of the limiting mixed Hodge structure through nearby-cycle hypercohomology. In this way, the present paper supplies the fully internal Hodge-theoretic refinement that was only identified as a gluing problem in [10].
The central point is that the corrected perverse object is not merely a constructible or perverse avatar of degeneration data, but admits a refinement internal to Saito’s category of mixed Hodge modules. This refinement is the missing link between the canonical perverse extension on the singular fiber and the Hodge-theoretic structure carried by nearby cycles. For a finite set of ordinary double points, the singular contribution is point-supported and rank one at each node, so the global problem becomes one of gluing the local ODP mixed-Hodge-module blocks into a single object on $X_{0}$ compatible with the variation morphism, the finite node-supported quotient, and the limiting mixed Hodge structure.
A second theme of the paper is that the finite multi-node setting carries nontrivial global extension-theoretic structure. The singular quotient is not just a bookkeeping device: it is a finite sum of localized rank-one sectors, one at each node, and the corresponding global corrected extension is organized by nodewise extension data. On the perverse side, this yields a distinguished finite-node extension picture which serves as the first precise mathematical shadow of the later quiver-, schober-, and wall-crossing directions. On the Hodge-theoretic side, the same quotient controls the finite vanishing sector in the limiting mixed Hodge structure. Thus the corrected extension is now visible simultaneously in $MHM(X_{0})$, in $\mathrm{Perv}(X_{0};\mathbb{Q})$, and on hypercohomology in the category of mixed Hodge structures.
1.1. Relation to earlier work
The starting point for the present paper is the construction carried out in [11]. There the corrected perverse object
$\mathcal{P}$
was defined in the ordinary double point case and shown to satisfy the structural properties expected of a canonical extension: it is perverse, restricts to the shifted constant sheaf on the smooth locus, is Verdier self-dual, and fits into an exact sequence
$0\to IC_{X_{0}}\to\mathcal{P}\to i_{*}\mathbb{Q}_{\{p\}}\to 0$
in the single-node case. Moreover, [10] established three points relevant here: (1) in §6 it proved that $\mathcal{P}$ is the unique minimal Verdier self-dual perverse extension of $\mathbb{Q}_{U}[3]$ across the node; (2) it framed the construction in terms of nearby and vanishing cycles, Picard–Lefschetz monodromy, and spherical-twist/schober-type categorical motivation; and (3) it studied the Hodge-theoretic content of the same perverse construction.
The main result in [10] was not the existence of a full mixed-Hodge-module lift, but rather the identification of a common nearby-cycle origin for the corrected perverse object and the corresponding degeneration data on the Hodge-theoretic side. In particular, we showed that the rank-one singular contribution in the perverse extension and the rank-one vanishing contribution in the limiting mixed Hodge structure arise from the same nearby-cycle and vanishing-cycle formalism, while isolating the explicit mixed-Hodge-module gluing problem as the next theorem to be proved.
The present paper fills precisely that gap. It is therefore neither a repetition of the perverse-sheaf arguments of [11] nor a mere reformulation of the Hodge-theoretic framework of [10]. Its role is to solve the missing construction problem: to build the mixed-Hodge-module refinement $\mathcal{P}^{H}$, prove that
$\mathrm{rat}(\mathcal{P}^{H})\cong\mathcal{P},$
establish the exact sequence in $MHM(X_{0})$, and identify the point-supported quotient with the finite vanishing part of the degeneration on the Hodge-theoretic side. In addition, the present paper makes explicit the finite-node extension-theoretic structure of the corrected object and its organization across the three levels $MHM$, perverse sheaves, and mixed Hodge structures.
1.2. Focused related work
The present paper draws on four closely related strands of work.
First, the sheaf-theoretic foundation comes from the formalism of perverse sheaves and recollement. The basic structural tools are those of Beilinson–Bernstein–Deligne [2], together with the linear-algebraic descriptions of perverse sheaves in the presence of isolated singularities due to MacPherson–Vilonen and Gelfand–MacPherson–Vilonen [7, 4]. These results provide the background for the canonical perverse extension on the singular fiber and its description in terms of open and closed gluing data.
Second, the local topological input comes from the classical theory of isolated hypersurface singularities. For an ordinary double point, the Milnor fiber has the homotopy type of a sphere in the middle degree, so the local vanishing cohomology is one-dimensional [8, 3]. This rank-one local structure is the source of the point-supported quotient appearing in the corrected perverse object and, on the Hodge side, of the local vanishing contribution to the nearby-cycle formalism.
Third, the Hodge-theoretic framework comes from Saito’s theory of mixed Hodge modules [13, 12]. In particular, Saito’s formalism provides:
  • the abelian categories $MHM(X)$,
  • the exact faithful realization functor
    $\mathrm{rat}:MHM(X)\to\mathrm{Perv}(X;\mathbb{Q}),$
  • nearby-cycle and vanishing-cycle functors in $MHM$,
  • and, crucially, the divisor-case gluing theorem for a principal divisor, expressed in terms of data $(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},u,v)$ satisfying the relation $vu=N$.
This divisor gluing formalism is the central technical mechanism used in the present paper. For expository background, we also make occasional use of Saito’s later overview [15].
Fourth, a methodological precedent for placing a perverse object built from nearby-cycle data into a mixed-Hodge-module framework is provided by the work of Banagl–Budur–Maxim [1]. Although their intersection-space complex is different from the canonical corrected perverse extension considered here, their paper is important because it shows, in a closely related isolated-singularity setting, how a perverse object constructed from nearby cycles can underlie a mixed Hodge module and thereby acquire canonical Hodge structures on hypercohomology. That work does not prove the present theorem, but it provides a useful model for the type of Hodge-theoretic refinement one should seek.
These references are the load-bearing ones for the present paper. Broader physical or categorical motivations, such as conifold transitions, perverse schobers, and spherical monodromy, are important for the larger program, but they are not the main technical input in the proofs below.
1.3. Physical and categorical motivation
Although the present paper is purely mathematical in its statements and proofs, the motivating geometry comes from the conifold transition picture in Calabi–Yau threefolds. In the ordinary double point case, the degeneration is governed by a collapsing three-sphere and the resulting rank-one Picard–Lefschetz monodromy on middle homology. In the physical interpretation of Strominger [19], this collapse corresponds to an additional light BPS state localized at the singularity. In the multi-node case, one expects a finite collection of such localized sectors, one for each node, together with global coupling data reflecting how the local contributions assemble into a single degeneration.
From the categorical side, the rank-one monodromy phenomena of an ODP are mirrored by the rank-one action of spherical twists and, more broadly, by perverse schober structures [17, 5]. The canonical corrected perverse object constructed in [11, 10] may therefore be regarded as the decategorified shadow of a local categorical monodromy phenomenon. The present paper adds the Hodge-theoretic layer needed to refine that picture: it identifies the mixed-Hodge-module object whose realization is the corrected perverse extension, whose quotient isolates the corresponding local vanishing sector on the Hodge-theoretic side, and whose global extension structure already exhibits the formal shape of a finite family of localized sectors coupled to a bulk geometric sector.
1.4. Main results
We now summarize the main theorem package of the paper.
Theorem 1.1 (Local mixed-Hodge-module ODP block).
Let
$\pi:\mathcal{X}\to\Delta$
be a one-parameter degeneration whose central fiber has a single ordinary double point $p$. Then there exists a local mixed-Hodge-module extension
$0\to IC^{H}_{\mathrm{loc}}\to\mathcal{P}^{H}_{\mathrm{loc}}\to i_{*}\mathbb{Q}^{H}_{\{p\}}(-1)\to 0$
whose realization is the local corrected perverse extension determined by nearby and vanishing cycles.
Theorem 1.2 (Finite multi-node support and extension structure).
Let
$\pi:\mathcal{X}\to\Delta$
be a one-parameter degeneration whose central fiber has ordinary double points
$\Sigma=\{p_{1},\dots,p_{r}\}.$
Then the singular quotient in the mixed-Hodge-module refinement is supported on $\Sigma$, and its point-supported contribution is canonically of the form
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Moreover, the corresponding global extension space decomposes nodewise, and on the perverse side the finite-node corrected extension admits a distinguished nodewise organization by local extension classes attached to the nodes.
Theorem 1.3 (Global gluing and realization).
There exists an object
$\mathcal{P}^{H}\in MHM(X_{0})$
fitting into an exact sequence
$0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\to 0$
such that
$\mathrm{rat}(\mathcal{P}^{H})\cong\mathcal{P}.$
Theorem 1.4 (Hypercohomology and the vanishing part of the LMHS).
The point-supported quotient
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
realizes the rank-$r$ local vanishing contribution in the nearby-cycle formalism. Moreover, the hypercohomology of $\mathcal{P}^{H}$ is functorially related to the limiting mixed Hodge structure through the same nearby-cycle mixed-Hodge-module construction, and the quotient contributes the vanishing part of the limiting mixed Hodge structure on hypercohomology.
Equivalently, the corrected extension is organized simultaneously at three levels: as a mixed-Hodge -module extension class, as its realized perverse extension class, and as an induced extension class in mixed Hodge structures on hypercohomology.
Taken together, these results provide the fully internal mixed-Hodge-module refinement of the canonical perverse extension in the finite multi-node ordinary double point case. They also make visible the nodewise extension-theoretic structure of the finite degeneration and furnish the precise Hodge-theoretic foundation needed for subsequent applications to global extension data, quiver-type structures, and the mathematical formulation of localized sectors in conifold-type physical models.
1.5. Proof strategy
The proof proceeds in four stages. We begin with the local ordinary double point model and construct the mixed-Hodge-module building block that refines the rank-one corrected perverse extension at a single node. This local step isolates the essential nearby-cycle, vanishing-cycle, and divisor-gluing data in Saito’s formalism and identifies the point-supported mixed Hodge module that refines the local vanishing contribution. The local construction is then used as the atomic input for the finite multi-node case.
The remaining stages pass from local to global. First, we prove that for finitely many ordinary double points the singular quotient is supported on the node set and decomposes into one rank-one point-supported mixed Hodge module per node. We then analyze the corresponding extension spaces and show that the finite-node corrected extension admits a nodewise organization. Second, we construct the global object $\mathcal{P}^{H}\in MHM(X_{0})$ by gluing the smooth-locus Hodge module to the finite collection of local node contributions and prove that its realization is the canonical corrected perverse object $\mathcal{P}$. Finally, we analyze the hypercohomology of $\mathcal{P}^{H}$ and show that the same nearby-cycle mixed-Hodge-module formalism governs the vanishing part of the limiting mixed Hodge structure. In this way, the local ODP model, the global extension on $X_{0}$, and the Hodge-theoretic degeneration data are placed in a single framework.
1.6. Scope and organization
The paper is confined to the case of finitely many ordinary double points. This is the natural next setting after the single-node theorem of [11] and the nearby-cycle bridge theorem of [10]. In particular, we do not attempt here a full treatment of arbitrary higher-dimensional singular strata, nor do we address in full generality the Kähler-package questions for the hypercohomology of the corrected object. Those problems remain downstream of the mixed-Hodge-module existence and realization theorem proved here.
Section 2 recalls the geometric setup of multi-node conifold degenerations, nearby and vanishing cycles, and the local topology of ordinary double points. Section 3 reviews the mixed-Hodge-module background needed later, with emphasis on nearby cycles, realization, and Saito’s divisor-case gluing theorem. Section 4 constructs the local mixed-Hodge-module ODP block. Section 5 proves the finite multi-node support decomposition and analyzes the corresponding extension structure. Section 6 gives the global gluing theorem and the realization theorem. Section 7 studies hypercohomology and proves the comparison with the vanishing part of the limiting mixed Hodge structure. Section 8 records auxiliary structural results and interprets the extension-theoretic output in quiver-type and domain-wall language. Section 9 discusses consequences and further directions.
2. Geometric setup and nearby-cycle background
Throughout the paper we work over the field $\mathbb{Q}$, and all varieties are complex algebraic varieties. Let
$\pi:\mathcal{X}\to\Delta$
be a one-parameter degeneration, where $\Delta$ is a small complex disk and $\Delta^{*}=\Delta\setminus\{0\}$. We write
$X_{t}:=\pi^{-1}(t)\qquad(t\in\Delta)$
for the fibers, and we assume that $X_{t}$ is smooth for $t\neq 0$. The central fiber is denoted
$X_{0}:=\pi^{-1}(0).$
Later, when projective Hodge-theoretic tools such as Hard Lefschetz are invoked, projectivity will be imposed explicitly. Until then, the constructions are formulated at the level of nearby cycles, perverse sheaves, and mixed Hodge modules for algebraic degenerations.
2.1. Conventions and normalizations
We use the middle perversity $t$-structure on the constructible derived category $D^{b}_{c}(X;\mathbb{Q})$, and write $\mathrm{Perv}(X;\mathbb{Q})$ for its heart [2, 6]. For a complex threefold, the shifted constant sheaf $\mathbb{Q}_{X}[3]$ is perverse on the smooth locus. Since the degenerations considered in this paper are degenerations of complex threefolds, we fix the normalization
$F:=\mathbb{Q}_{\mathcal{X}}[3].$
With this convention, the nearby-cycle and vanishing-cycle objects attached to $F$ lie in the perverse heart after the standard normalization of the functors $^{p}\psi_{\pi}$ and $^{p}\phi_{\pi}$ [3, Chapter 6].
To simplify notation, we suppress the superscript p and write
$\psi_{\pi}(F),\qquad\phi_{\pi}(F)$
for the perverse nearby-cycle and vanishing-cycle objects. Likewise, we write
$\mathrm{can}_{F}:\psi_{\pi}(F)\to\phi_{\pi}(F),\qquad\mathrm{var}_{F}:\phi_{\pi}(F)\to\psi_{\pi}(F)$
for the canonical and variation morphisms. These satisfy the usual relation
$\mathrm{can}_{F}\circ\mathrm{var}_{F}=T-\operatorname{id},\qquad\mathrm{var}_{F}\circ\mathrm{can}_{F}=T-\operatorname{id},$
where $T$ denotes the local monodromy operator on nearby cycles [3, Chapter 6]. All cones appearing below are taken in the derived category $D^{b}_{c}(X_{0};\mathbb{Q})$.
2.2. Multi-node conifold degenerations
We now specialize to the case in which the central fiber $X_{0}$ has only finitely many ordinary double points
$\Sigma=\{p_{1},\dots,p_{r}\}\subset X_{0}.$
We refer to such a degeneration as a finite multi-node conifold degeneration. Write
$U:=X_{0}\setminus\Sigma$
for the smooth locus of the central fiber, and let
$j:U\hookrightarrow X_{0},\qquad i_{k}:\{p_{k}\}\hookrightarrow X_{0}$
denote the natural inclusions.
The point of this setup is that the singular locus is zero-dimensional, so the singular contribution to the nearby-cycle formalism is concentrated at finitely many isolated points. In particular, the corrected object we shall construct differs from the intersection complex of $X_{0}$ only by a finite point-supported contribution. [11] establishes this in the single-node case, while [10] isolates the corresponding multi-node structure at the level of perverse sheaves and nearby-cycle/Hodge-theoretic compatibility [11, 10].
2.3. Local topology of an ordinary double point
Let $p\in X_{0}$ be an ordinary double point. Then locally near $p$, the singularity is an isolated hypersurface singularity, and its Milnor fiber $F_{p}$ has the homotopy type of a sphere of real dimension $3$ in the threefold case [8, 3]. Hence
$\widetilde{H}^{k}(F_{p};\mathbb{Q})\cong\begin{cases}\mathbb{Q}&k=3,\\ 0&k\neq 3.\end{cases}$
Equivalently, the local vanishing cohomology is one-dimensional and concentrated in the middle degree.
It follows that the vanishing-cycle contribution of an ordinary double point is rank one. After the perverse normalization adopted above, this gives a point-supported perverse object whose local stalk at the node is one-dimensional. In the finite multi-node case, each node therefore contributes one rank-one local vanishing summand. This is the local topological source of the singular quotient that appears later in both the perverse-sheaf and mixed-Hodge-module exact sequences.
2.4. Nearby cycles, vanishing cycles, and variation
Let
$F:=\mathbb{Q}_{\mathcal{X}}[3].$
The nearby-cycle and vanishing-cycle functors fit into the standard distinguished triangles in $D^{b}_{c}(X_{0};\mathbb{Q})$; see, for example, [3, Chapter 6]. In particular, there are functorial morphisms
$\mathrm{can}_{F}:\psi_{\pi}(F)\to\phi_{\pi}(F),\qquad\mathrm{var}_{F}:\phi_{\pi}(F)\to\psi_{\pi}(F),$
which encode the relation between the specialization to the singular fiber and the local vanishing contribution.
In the threefold ordinary double point case, the object $\phi_{\pi}(F)$ is supported on the finite set $\Sigma$, and each local contribution is one-dimensional by the Milnor-fiber calculation above. Thus the nontrivial singular contribution of the degeneration enters entirely through the variation morphism between the vanishing-cycle and nearby-cycle objects.
This is the formal setting in which the corrected perverse object is defined.
2.5. The corrected perverse object in the multi-node setting
We define the corrected object by
$\mathcal{P}:=\operatorname{Cone}(\mathrm{var}_{F}:\phi_{\pi}(F)\to\psi_{\pi}(F))[-1].$
This is the multi-node analogue of the canonical object constructed in the single-node case in [11]. In that paper, the single-node ordinary double point case was treated in detail: the corresponding object was shown to be perverse and to fit into an exact sequence
$0\to IC_{X_{0}}\to\mathcal{P}\to i_{*}\mathbb{Q}_{\{p\}}\to 0.$
In [10], $\mathcal{P}$ was shown to be Verdier self-dual and the construction was placed in the framework of nearby-cycle formalism and limiting mixed Hodge theory, and the existence of a fully internal mixed-Hodge-module refinement was identified as the main open gluing problem.
The purpose of the present paper is to solve that gluing problem in the finite multi-node case. Thus $\mathcal{P}$ is not merely an auxiliary perverse object: it is the constructible/perverse shadow of the mixed-Hodge-module object that we shall construct later.
2.6. Picard–Lefschetz data and the vanishing lattice
For each node $p_{k}\in\Sigma$, let
$\delta_{k}\in H_{3}(X_{t},\mathbb{Z})$
denote a corresponding local vanishing cycle. The local monodromy about $t=0$ is governed by the Picard–Lefschetz formula, and the span of the $\delta_{k}$ in middle homology carries the vanishing-cycle lattice of the degeneration [8, 3]. In particular, the collection
$\delta_{1},\dots,\delta_{r}$
determines a rank-$r$ local vanishing contribution together with the intersection pairing among the vanishing cycles.
For the present paper, we use this data only at the level of geometric motivation and support decomposition: each $\delta_{k}$ gives one point-supported rank-one contribution on the singular fiber. The more refined question of how the intersection relations among the vanishing cycles are reflected in the global extension data will reappear later when we discuss the global gluing problem and its possible quiver-theoretic shadow.
3. Mixed Hodge modules and divisor gluing
The purpose of this section is to record the mixed-Hodge-module formalism that will be used in the construction of the global refinement
$\mathcal{P}^{H}\in MHM(X_{0}).$
We restrict attention to the exact properties of Saito’s theory that are needed later: the existence of the abelian categories $MHM(X)$, the realization functor, nearby and vanishing cycles in the mixed-Hodge-module setting, and the principal-divisor gluing formalism. These are the ingredients already isolated in [10] as the formal input required for a full Hodge-theoretic refinement of the corrected perverse extension.
3.1. Mixed Hodge modules and realization
Let $X$ be a complex algebraic variety. Saito constructs an abelian category $MHM(X)$ of mixed Hodge modules on $X$, together with an exact and faithful functor
$\mathrm{rat}:MHM(X)\to\mathrm{Perv}(X;\mathbb{Q}),$
to the category of rational perverse sheaves on $X$ [13, 12]. We refer to $\mathrm{rat}$ as the realization functor. In particular, every mixed Hodge module has an underlying rational perverse sheaf, and exact sequences in $MHM(X)$ remain exact after applying $\mathrm{rat}$.
For smooth varieties, Saito’s theory extends the classical category of admissible variations of mixed Hodge structure. More generally, it provides a Hodge-theoretic enhancement of the formalism of perverse sheaves and regular holonomic $D$-modules compatible with the standard functorial operations. For the present paper, the key point is that the category $MHM(X_{0})$ is rich enough to support both the intersection complex
$IC^{H}_{X_{0}}$
and the point-supported mixed Hodge modules
$i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
that will appear as the singular quotient in the corrected extension.
3.2. Nearby and vanishing cycles in MHMMHM
Let $\pi:\mathcal{X}\to\Delta$ be a one-parameter degeneration, and let $X_{0}=\pi^{-1}(0)$. Saito defines nearby-cycle and vanishing-cycle functors in the mixed-Hodge-module setting,
$\psi_{\pi}^{H},\qquad\phi_{\pi}^{H},$
which are compatible with the corresponding perverse nearby-cycle and vanishing-cycle functors under the realization functor [13, 12]. Thus, for every object $\mathcal{M}\in MHM(\mathcal{X})$, one has
$\mathrm{rat}\!\bigl(\psi_{\pi}^{H}(\mathcal{M})\bigr)\cong\psi_{\pi}\!\bigl(\mathrm{rat}(\mathcal{M})\bigr),\qquad\mathrm{rat}\!\bigl(\phi_{\pi}^{H}(\mathcal{M})\bigr)\cong\phi_{\pi}\!\bigl(\mathrm{rat}(\mathcal{M})\bigr),$
with the usual normalized perverse nearby-cycle and vanishing-cycle functors on the right-hand side.
These functors carry canonical and variation morphisms,
$\mathrm{can}^{H}:\psi_{\pi}^{H}(\mathcal{M})\to\phi_{\pi}^{H}(\mathcal{M}),\qquad\mathrm{var}^{H}:\phi_{\pi}^{H}(\mathcal{M})\to\psi_{\pi}^{H}(\mathcal{M}),$
whose realizations are the corresponding canonical and variation morphisms of perverse sheaves. Moreover, the nilpotent monodromy operator
$N:=\log T_{u}$
acts in the mixed-Hodge-module formalism and controls the relation between nearby and vanishing cycles. This is the Hodge-theoretic refinement of the same nearby-cycle/vanishing-cycle mechanism that defines the corrected perverse object
$\mathcal{P}=\operatorname{Cone}(\mathrm{var}_{F})[-1]$
on the perverse side.
In particular, if
$F=\mathbb{Q}_{\mathcal{X}}[3],$
then the nearby-cycle and vanishing-cycle mixed Hodge modules attached to $F$ refine the perverse objects $\psi_{\pi}(F)$ and $\phi_{\pi}(F)$, and the realization functor carries the mixed-Hodge-module variation morphism $\mathrm{var}_{F}^{H}$ to the perverse variation morphism $\mathrm{var}_{F}$.
3.3. Saito’s divisor gluing theorem
The formal engine of the present paper is Saito’s gluing theorem for a principal divisor [13, Prop. 0.3]. Let $g:X\to\mathbb{C}$ be a regular function, and let
$Y:=g^{-1}(0),\qquad U:=X\setminus Y.$
Then mixed Hodge modules on $X$ whose behavior is controlled along the divisor $Y$ may be described in terms of gluing data consisting of:
  • an object on the complement $U$,
  • an object on the divisor $Y$,
  • and morphisms
    $u:\psi_{g,1}(\mathcal{M}^{\prime})\to\mathcal{M}^{\prime\prime},\qquad v:\mathcal{M}^{\prime\prime}\to\psi_{g,1}(\mathcal{M}^{\prime})(-1),$
    satisfying the relation
    $vu=N,$
    where $N$ is the nilpotent monodromy operator.
Here $\psi_{g,1}$ denotes the unipotent nearby-cycle part.
In the present situation, the central fiber
$X_{0}=\pi^{-1}(0)$
is a principal divisor in $\mathcal{X}$. Consequently, the problem of constructing a mixed-Hodge-module refinement of the corrected perverse object is reduced to the explicit identification of the gluing datum
$(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},u,v)$
whose realization recovers the variation morphism
$\mathrm{var}_{F}:\phi_{\pi}(F)\to\psi_{\pi}(F)$
and hence the cone object
$\mathcal{P}=\operatorname{Cone}(\mathrm{var}_{F})[-1].$
This is precisely the gluing problem isolated in [10] in the single-node case, and the present paper solves it in the finite multi-node ordinary double point setting.
3.4. Point-supported mixed Hodge modules
For each node $p_{k}\in\Sigma$, let
$i_{k}:\{p_{k}\}\hookrightarrow X_{0}$
denote the inclusion. The point $\{p_{k}\}$ supports the pure Hodge module
$\mathbb{Q}^{H}_{\{p_{k}\}},$
whose realization is the skyscraper perverse sheaf $\mathbb{Q}_{\{p_{k}\}}$. We write
$i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
for its Tate-twisted pushforward to $X_{0}$. Its realization is
$\mathrm{rat}\!\bigl(i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\bigr)\cong i_{k*}\mathbb{Q}_{\{p_{k}\}},$
since the Tate twist modifies the Hodge-theoretic structure but does not alter the underlying rational perverse sheaf.
These point-supported mixed Hodge modules are the natural Hodge-theoretic refinements of the rank-one local singular contributions appearing in the corrected perverse extension. In the single-node case, [11] and [10] identify the quotient
$i_{*}\mathbb{Q}_{\{p\}}$
as the point-supported rank-one perverse correction, and the corresponding mixed-Hodge-module object
$i_{*}\mathbb{Q}^{H}_{\{p\}}(-1)$
was isolated there as the expected singular quotient of the missing Hodge-theoretic refinement. The multi-node theorem of the present paper will show that in the finite-node case the singular quotient is precisely
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
3.5. Tate twists, weights, and normalization conventions
We now fix the Hodge-theoretic normalization used throughout the paper. We follow Saito’s Tate twist conventions: for a mixed Hodge module $\mathcal{M}$, the twist $\mathcal{M}(-1)$ lowers weights by $2$, and the nilpotent monodromy operator satisfies
$N:\psi_{\pi,1}^{H}(\mathcal{M})\longrightarrow\psi_{\pi,1}^{H}(\mathcal{M})(-1).$
Accordingly, the point-supported rank-one singular term refining a local vanishing contribution is written
$i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),$
not as a half-integral Tate object. This convention is essential: all Hodge-theoretic statements in the paper are formulated inside the standard category of rational mixed Hodge modules and rational mixed Hodge structures, where Tate twists are indexed by integers.
In particular, when we speak of the rank-one local vanishing contribution of an ordinary double point, we mean the point-supported mixed-Hodge-module summand
$i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
and its contribution under hypercohomology to the limiting mixed Hodge structure. The present paper does not use any fractional Tate twist notation. All later exact sequences, support decompositions, and hypercohomology comparisons are written with this normalization.
This choice is compatible with the role of the monodromy operator in Saito’s formalism and with the point-supported quotient anticipated in [10]. It is also the normalization needed for the global exact sequence in $MHM(X_{0})$ proved later in the paper.
4. Local ODP construction via nearby-cycle quotients
In this section we construct the local mixed-Hodge-module block attached to an ordinary double point by an explicit quotient procedure in the nearby/vanishing-cycle formalism. The key point is that the desired local point-supported object is not introduced abstractly by a lifting ansatz, but is extracted canonically from the $\mathrm{can}/\mathrm{Var}/N$ package together with the one-dimensional vanishing sector of the ordinary double point.
4.1. Local analytic model and notation
Let
$\pi_{\mathrm{loc}}:\mathcal{X}_{\mathrm{loc}}\to\Delta$
be a local one-parameter degeneration whose central fiber
$X_{0,\mathrm{loc}}:=\pi_{\mathrm{loc}}^{-1}(0)$
has a single ordinary double point at
$p\in X_{0,\mathrm{loc}}.$
Write
$U_{\mathrm{loc}}:=X_{0,\mathrm{loc}}\setminus\{p\},\qquad j_{\mathrm{loc}}:U_{\mathrm{loc}}\hookrightarrow X_{0,\mathrm{loc}},\qquad i_{\mathrm{loc}}:\{p\}\hookrightarrow X_{0,\mathrm{loc}}.$
Let
$F_{\mathrm{loc}}:=\mathbb{Q}_{\mathcal{X}_{\mathrm{loc}}}[3].$
We work with the unipotent nearby- and vanishing-cycle functors in Saito’s category of mixed Hodge modules:
$\psi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}}),\qquad\phi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}}),$
together with the canonical maps
$\mathrm{can}^{H}:\psi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})\longrightarrow\phi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}}),\qquad\mathrm{Var}^{H}:\phi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})\longrightarrow\psi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})(-1),$
satisfying
$\mathrm{Var}^{H}\circ\mathrm{can}^{H}=N$
on the unipotent nearby-cycle object [14].
4.2. The local point-supported target
The Milnor fiber of an ordinary double point has one-dimensional vanishing cohomology in middle degree. Equivalently, the local vanishing sector is a one-dimensional mixed Hodge structure of Tate type [9]. Accordingly, we define the point-supported local mixed Hodge module
$W^{H}_{\mathrm{loc}}:=i_{\mathrm{loc}*}\mathbb{Q}^{H}_{\{p\}}(-1).$
Its underlying rational perverse sheaf is the point-supported rank-one perverse block
$K_{\mathrm{ODP}}:=\mathrm{rat}\bigl(W^{H}_{\mathrm{loc}}\bigr).$
Lemma 4.1 (Local ODP zig-zag block).
The point-supported local perverse block $K_{\mathrm{ODP}}$ has MacPherson–Vilonen zig-zag
$Z(K_{\mathrm{ODP}})\cong(0,\mathbb{Q},\mathbb{Q},0,\operatorname{id},0).$
In particular, the corresponding map
$\beta:H^{0}(i_{\mathrm{loc}}^{!}K_{\mathrm{ODP}})\longrightarrow H^{0}(i_{\mathrm{loc}}^{*}K_{\mathrm{ODP}})$
is an isomorphism, and therefore
$\mathrm{End}_{\mathrm{Perv}(X_{0,\mathrm{loc}};\mathbb{Q})}(K_{\mathrm{ODP}})\cong\mathbb{Q}.$
Proof.
Since $K_{\mathrm{ODP}}$ is point-supported, one has
$j_{\mathrm{loc}}^{*}K_{\mathrm{ODP}}=0.$
Hence
$i_{\mathrm{loc}}^{*}Rj_{\mathrm{loc}*}j_{\mathrm{loc}}^{*}K_{\mathrm{ODP}}=0,$
and the MacPherson–Vilonen zig-zag exact sequence reduces to
$0\longrightarrow H^{0}(i_{\mathrm{loc}}^{!}K_{\mathrm{ODP}})\xrightarrow{\beta}H^{0}(i_{\mathrm{loc}}^{*}K_{\mathrm{ODP}})\longrightarrow 0.$
Thus $\beta$ is an isomorphism. Since both middle terms are one-dimensional, the zig-zag is
$(0,\mathbb{Q},\mathbb{Q},0,\operatorname{id},0).$
MacPherson–Vilonen then imply that the local Hom-space is computed exactly by zig-zags in this situation, so the endomorphism ring is one-dimensional [7]. ∎
Lemma 4.1 rigidifies the rational perverse target: any nonzero morphism to or from $K_{\mathrm{ODP}}$ is unique up to scalar.
4.3. The local quotient object
The ambient nearby/vanishing-cycle package contains more information than the local ODP correction block. The correct local quotient is extracted from the $\mathrm{can}/\mathrm{Var}$ interface.
Definition 4.2.
Define
$A_{\mathrm{loc}}:=\operatorname{Im}(\mathrm{can}^{H})\cap\ker(\mathrm{Var}^{H})\subset\phi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}}).$
This is the local object that survives both the incoming nearby-cycle map and the vanishing under the outgoing variation map. It is the natural local analogue of the monodromy-controlled quotient constructions that appear in the isolated-singularity setting [1].
Proposition 4.3 (Point-level ODP quotient).
The point-level mixed Hodge structure
$H^{0}\bigl(i_{\mathrm{loc}}^{*}A_{\mathrm{loc}}\bigr)$
admits a nonzero quotient
$q_{\mathrm{pt}}:H^{0}\bigl(i_{\mathrm{loc}}^{*}A_{\mathrm{loc}}\bigr)\twoheadrightarrow\mathbb{Q}^{H}(-1)$
corresponding to the one-dimensional ordinary-double-point vanishing line. Consequently, by adjunction for the closed immersion $i_{\mathrm{loc}}$, there is an induced nonzero morphism
$q_{\mathrm{loc}}:A_{\mathrm{loc}}\longrightarrow i_{\mathrm{loc}*}\mathbb{Q}^{H}_{\{p\}}(-1)=W^{H}_{\mathrm{loc}}.$
Proof.
For an ordinary double point, the Milnor fiber has one-dimensional vanishing cohomology in middle degree. The local nearby/vanishing-cycle formalism therefore determines a one-dimensional vanishing mixed Hodge structure of Tate type $\mathbb{Q}^{H}(-1)$ on the local ODP sector. Since
$A_{\mathrm{loc}}=\operatorname{Im}(\mathrm{can}^{H})\cap\ker(\mathrm{Var}^{H}),$
its point-level fiber
$H^{0}\bigl(i_{\mathrm{loc}}^{*}A_{\mathrm{loc}}\bigr)$
admits a nonzero quotient onto that one-dimensional vanishing line. This defines
$q_{\mathrm{pt}}:H^{0}\bigl(i_{\mathrm{loc}}^{*}A_{\mathrm{loc}}\bigr)\twoheadrightarrow\mathbb{Q}^{H}(-1).$
Adjunction for $i_{\mathrm{loc}}$ then yields the induced morphism
$q_{\mathrm{loc}}:A_{\mathrm{loc}}\to i_{\mathrm{loc}*}\mathbb{Q}^{H}_{\{p\}}(-1).$
Since $q_{\mathrm{pt}}$ is nonzero, so is $q_{\mathrm{loc}}$. ∎
Remark 4.1.
The quotient $q_{\mathrm{pt}}$ is canonical only relative to the distinguished one-dimensional ordinary-double-point vanishing line singled out by the local nearby-/vanishing-cycle formalism. Thus the construction is canonical up to the standard scalar normalization already inherent in the rank-one ODP sector.
Definition 4.4.
Define
$u_{\mathrm{loc}}:=q_{\mathrm{loc}}\circ\mathrm{can}^{H}:\psi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})\longrightarrow W^{H}_{\mathrm{loc}}.$
Proposition 4.5 (Nonvanishing of the realized local map).
The realized morphism
$\mathrm{rat}(u_{\mathrm{loc}}):\psi_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})\longrightarrow K_{\mathrm{ODP}}$
is nonzero.
Proof.
By Proposition 4.3, the morphism
$q_{\mathrm{loc}}:A_{\mathrm{loc}}\to W^{H}_{\mathrm{loc}}$
is nonzero and is induced from the ordinary-double-point vanishing line. Since $\mathrm{can}^{H}$ is the canonical nearby-to-vanishing morphism in the unipotent local formalism, its restriction to the local ODP sector is nontrivial. Therefore
$u_{\mathrm{loc}}=q_{\mathrm{loc}}\circ\mathrm{can}^{H}$
is nonzero on that sector. Exactness of realization then implies that
$\mathrm{rat}(u_{\mathrm{loc}})$
is nonzero. ∎
Proposition 4.6 (Uniqueness of $u_{\mathrm{loc}}$).
The morphism $u_{\mathrm{loc}}$ is unique up to nonzero scalar.
Proof.
Consider the realization map
$\mathrm{rat}:\mathrm{Hom}_{MHM(X_{0,\mathrm{loc}})}\bigl(\psi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}}),W^{H}_{\mathrm{loc}}\bigr)\longrightarrow\mathrm{Hom}_{\mathrm{Perv}(X_{0,\mathrm{loc}};\mathbb{Q})}\bigl(\psi_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}}),K_{\mathrm{ODP}}\bigr).$
Realization is faithful, hence injective on Hom-sets. The target is one-dimensional by the local ODP zig-zag rigidity. Since $\mathrm{rat}(u_{\mathrm{loc}})\neq 0$ by Proposition 4.5, it spans the target. Therefore $u_{\mathrm{loc}}$ is unique up to nonzero scalar. ∎
4.4. The dual local map
The dual local map is constructed symmetrically from the same point-supported block.
Definition 4.7.
Let
$v_{\mathrm{loc}}:W^{H}_{\mathrm{loc}}\longrightarrow\psi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})(-1)$
denote the unique nonzero morphism induced from $\mathrm{Var}^{H}$ on the local ODP sector.
Proposition 4.8 (Existence and uniqueness of the dual local map).
There exists a nonzero morphism
$v_{\mathrm{loc}}:W^{H}_{\mathrm{loc}}\longrightarrow\psi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})(-1)$
induced from $\mathrm{Var}^{H}$ on the local ODP sector. It is unique up to nonzero scalar.
Proof.
The ordinary-double-point local sector is one-dimensional, and the point-supported target block
$W^{H}_{\mathrm{loc}}=i_{\mathrm{loc}*}\mathbb{Q}^{H}_{\{p\}}(-1)$
is uniquely determined by that sector. The variation morphism
$\mathrm{Var}^{H}:\phi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})\longrightarrow\psi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})(-1)$
is nontrivial on the same local ODP vanishing line, hence induces a nonzero morphism
$v_{\mathrm{loc}}:W^{H}_{\mathrm{loc}}\to\psi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})(-1).$
After realization, this yields a nonzero morphism from the rigid perverse block $K_{\mathrm{ODP}}$. Since the corresponding perverse Hom-space is one-dimensional, uniqueness up to scalar follows. Faithfulness of realization then implies the same uniqueness upstairs in $MHM$. ∎
4.5. Normalization by monodromy
Proposition 4.9 (Local normalization).
After rescaling $u_{\mathrm{loc}}$ and $v_{\mathrm{loc}}$, one may arrange that
$v_{\mathrm{loc}}u_{\mathrm{loc}}=N$
on the local ordinary-double-point sector.
Proof.
In Saito’s local formalism one has
$\mathrm{Var}^{H}\circ\mathrm{can}^{H}=N.$
By Propositions 4.6 and 4.8, both $u_{\mathrm{loc}}$ and $v_{\mathrm{loc}}$ are nonzero and unique up to scalar on the one-dimensional ODP sector. Hence the composite $v_{\mathrm{loc}}u_{\mathrm{loc}}$ differs from the induced action of NN on that sector by a nonzero scalar. Rescaling one of the two maps removes that scalar. ∎
4.6. Local gluing datum
We now have the local data required for Saito’s gluing formalism. At this point all hypotheses of Saito’s local divisor gluing formalism are verified on the ODP sector.
Theorem 4.10 (Local ODP gluing datum).
The quadruple
$\bigl(\mathbb{Q}^{H}_{U_{\mathrm{loc}}}[3],\,W^{H}_{\mathrm{loc}},\,u_{\mathrm{loc}},\,v_{\mathrm{loc}}\bigr)$
is a local gluing datum on the ordinary-double-point sector, with
$W^{H}_{\mathrm{loc}}=i_{\mathrm{loc}*}\mathbb{Q}^{H}_{\{p\}}(-1),\qquad v_{\mathrm{loc}}u_{\mathrm{loc}}=N.$
Proof.
This follows immediately from Definition 4.4 and Propositions 4.8, 4.9. ∎
4.7. The local mixed-Hodge-module correction block
Theorem 4.11 (Local ODP mixed-Hodge-module extension).
There exists a local mixed Hodge module
$\mathcal{P}^{H}_{\mathrm{loc}}$
fitting into an exact sequence
$0\to IC^{H}_{\mathrm{loc}}\to\mathcal{P}^{H}_{\mathrm{loc}}\to W^{H}_{\mathrm{loc}}\to 0,$
where
$W^{H}_{\mathrm{loc}}=i_{\mathrm{loc}*}\mathbb{Q}^{H}_{\{p\}}(-1).$
Proof.
Apply Saito’s local gluing formalism to the datum of Theorem 4.10. The hypotheses are satisfied by Theorem 4.10, with bulk term $IC^{H}_{\mathrm{loc}}$ and singular term $W^{H}_{\mathrm{loc}}$, so the gluing theorem produces the required extension object. ∎
4.8. Realization and the local corrected perverse object
Theorem 4.12 (Local realization).
The realization of $\mathcal{P}^{H}_{\mathrm{loc}}$ is the corrected local perverse object:
$\mathrm{rat}(\mathcal{P}^{H}_{\mathrm{loc}})\cong\mathcal{P}_{\mathrm{loc}}.$
Proof.
The realized local datum is
$\bigl(\mathbb{Q}_{U_{\mathrm{loc}}}[3],\,K_{\mathrm{ODP}},\,\mathrm{rat}(u_{\mathrm{loc}}),\,\mathrm{rat}(v_{\mathrm{loc}})\bigr),$
where $K_{\mathrm{ODP}}$ is the rigid local point-supported perverse block and $\mathrm{rat}(u_{\mathrm{loc}})$, $\mathrm{rat}(v_{\mathrm{loc}})$ are the unique nonzero local perverse maps by Propositions 4.5, 4.6, 4.8. Since realization is exact and carries the mixed-Hodge-module gluing maps to the corresponding perverse gluing maps, the realized extension is precisely the corrected local perverse ODP extension. ∎
4.9. Cone description
Proposition 4.13 (Local cone description).
In the derived category of mixed Hodge modules, one has
$\mathcal{P}^{H}_{\mathrm{loc}}\simeq\operatorname{Cone}(v_{\mathrm{loc}})[-1].$
Proof.
The local gluing datum determines a distinguished triangle
$W^{H}_{\mathrm{loc}}\longrightarrow\psi^{H}_{\pi_{\mathrm{loc}},1}(F_{\mathrm{loc}})(-1)\longrightarrow\mathcal{P}^{H}_{\mathrm{loc}}[1]\overset{+1}{\longrightarrow}.$
By definition of cone in the triangulated category, the third term is
$\operatorname{Cone}(v_{\mathrm{loc}}),$
hence
$\mathcal{P}^{H}_{\mathrm{loc}}\simeq\operatorname{Cone}(v_{\mathrm{loc}})[-1].$
4.10. Local rigidity
Proposition 4.14 (Local rigidity).
Any local mixed Hodge module $\mathcal{E}^{H}$ whose realization is the corrected local perverse ODP block and whose singular quotient is $W^{H}_{\mathrm{loc}}$ is isomorphic to $\mathcal{P}^{H}_{\mathrm{loc}}$.
Proof.
The realized local datum is rigid by Lemma 4.1, and the corresponding mixed-Hodge-module maps are unique up to scalar by Propositions 4.6 and 4.8. Therefore the gluing datum is unique up to the same scalar normalizations already fixed by Proposition 4.9. Hence
$\mathcal{E}^{H}\cong\mathcal{P}^{H}_{\mathrm{loc}}.$
5. Finite multi-node support decomposition
We now pass from the local ordinary double point block to the finite multi-node setting. Let
$\Sigma=\{p_{1},\dots,p_{r}\}\subset X_{0}$
be the set of ordinary double points of the central fiber, and write
$U:=X_{0}\setminus\Sigma,\qquad j:U\hookrightarrow X_{0},\qquad i_{k}:\{p_{k}\}\hookrightarrow X_{0}.$
The purpose of this section is to identify the singular contribution relevant to the global mixed-Hodge-module refinement as a finite direct sum of the local rank-one point-supported blocks constructed in Section 4, and to describe the corresponding global extension space in nodewise form. This is the first genuine local-to-global step in the paper: the local ODP mixed-Hodge-module building block is now assembled over the finite singular set.
5.1. Support of the vanishing-cycle object
Let
$F:=\mathbb{Q}_{\mathcal{X}}[3].$
Since the central fiber $X_{0}$ has only isolated ordinary double points, the singular locus of the degeneration is the finite set $\Sigma$. It follows from the constructibility of nearby and vanishing cycles that the vanishing-cycle object
$\phi_{\pi}^{H}(F)$
is supported on $\Sigma$. Equivalently, away from the finite node set there is no local vanishing contribution.
Lemma 5.1.
The support of the vanishing-cycle mixed Hodge module $\phi_{\pi}^{H}(F)$ is contained in the finite set $\Sigma$.
Proof.
Vanishing cycles detect the failure of local topological triviality of the degeneration. Since the fibers are smooth away from the nodes of the central fiber, the only possible nontrivial vanishing contribution occurs at the points $p_{k}\in\Sigma$. Therefore
$\operatorname{supp}\bigl(\phi_{\pi}^{H}(F)\bigr)\subseteq\Sigma.$
The same statement holds on the perverse side, and compatibility with realization gives the mixed-Hodge-module version. ∎
5.2. The finite direct sum of local ODP blocks
Section 4 constructs, at each node $p_{k}$, a point-supported local mixed-Hodge-module block
$W_{k}^{H}:=i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),$
whose realization is the rigid local ODP perverse block at $p_{k}$. These local blocks are supported on pairwise disjoint closed points, so they assemble canonically into a finite point-supported mixed Hodge module
$Q_{\Sigma}^{H}:=\bigoplus_{k=1}^{r}W_{k}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Proposition 5.2.
There is a canonically defined finite point-supported mixed Hodge module
$Q_{\Sigma}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
obtained by assembling the local ODP quotient blocks of Section 4 over the finite node set $\Sigma$.
Proof.
For each node $p_{k}$, Section 4 constructs the local point-supported block
$W_{k}^{H}=i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Since the supports $\{p_{k}\}$ are pairwise disjoint, these objects may be summed in $MHM(X_{0})$, giving the finite point-supported object
$Q_{\Sigma}^{H}:=\bigoplus_{k=1}^{r}W_{k}^{H}.$
This construction is canonical once the local ODP blocks are fixed. ∎
5.3. Node-wise singular quotient theorem
The role of the preceding proposition is not to identify the whole vanishing-cycle mixed Hodge module, but rather to identify the singular quotient object that enters the global corrected extension.
Proposition 5.3.
The singular quotient in the finite multi-node mixed-Hodge-module refinement is
$Q_{\Sigma}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Proof.
At each node $p_{k}$, the local ordinary double point theorem of Section 4 identifies the singular quotient block as
$W_{k}^{H}=i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
The finite node set $\Sigma$ is a disjoint union of these closed points, so the total singular quotient object appearing in the global corrected extension is the direct sum of the local quotient blocks:
$Q_{\Sigma}^{H}=\bigoplus_{k=1}^{r}W_{k}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Remark 5.1.
Proposition 5.3 does not assert that the entire vanishing-cycle mixed Hodge module $\phi_{\pi}^{H}(F)$ splits as the direct sum of the node blocks. Rather, it identifies the finite point-supported singular quotient object relevant to the corrected extension constructed from the local ODP sectors.
5.4. Nodewise decomposition of the extension space
The global corrected extension is classified by an extension class whose source is the finite point-supported singular quotient object $Q_{\Sigma}^{H}$ and whose target is the bulk intersection-complex Hodge module $IC^{H}_{X_{0}}$. We first record the formal additivity of the relevant Ext-groups over the finite node set.
Lemma 5.4.
In the abelian category $\mathrm{Perv}(X_{0};\mathbb{Q})$, one has a natural isomorphism
$\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\Bigl(\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}},IC_{X_{0}}\Bigr)\cong\bigoplus_{k=1}^{r}\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\bigl(i_{k*}\mathbb{Q}_{\{p_{k}\}},IC_{X_{0}}\bigr).$
Proof.
The functor $\mathrm{Hom}_{\mathrm{Perv}(X_{0};\mathbb{Q})}(-,IC_{X_{0}})$ sends finite direct sums in the first variable to finite direct products, and for a finite index set these products coincide with direct sums. Passing to the first right-derived functor yields the stated decomposition of $\mathrm{Ext}^{1}$. ∎
Lemma 5.5.
In the abelian category $MHM(X_{0})$, one has a natural isomorphism
$\mathrm{Ext}^{1}_{MHM(X_{0})}\bigl(Q_{\Sigma}^{H},IC^{H}_{X_{0}}\bigr)\cong\bigoplus_{k=1}^{r}\mathrm{Ext}^{1}_{MHM(X_{0})}\bigl(i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),IC^{H}_{X_{0}}\bigr).$
Proof.
The same finite-additivity argument applies in the abelian category $MHM(X_{0})$, using the finite direct-sum decomposition
$Q_{\Sigma}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Applying the first derived functor of $\mathrm{Hom}_{MHM(X_{0})}(-,IC^{H}_{X_{0}})$ gives the result. ∎
Proposition 5.6.
The global corrected extension class lies in
$\mathrm{Ext}^{1}_{MHM(X_{0})}\bigl(Q_{\Sigma}^{H},IC^{H}_{X_{0}}\bigr),$
and under the decomposition of Lemma 5.5 it determines, and is equivalently determined by, a tuple of nodewise extension classes
$(\epsilon_{1}^{H},\dots,\epsilon_{r}^{H}),\qquad\epsilon_{k}^{H}\in\mathrm{Ext}^{1}_{MHM(X_{0})}\bigl(i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),IC^{H}_{X_{0}}\bigr).$
Proof.
By Proposition 5.3, the singular quotient object in the global corrected extension is $Q_{\Sigma}^{H}$. Any such global corrected extension therefore defines a class in
$\mathrm{Ext}^{1}_{MHM(X_{0})}(Q_{\Sigma}^{H},IC^{H}_{X_{0}}).$
Lemma 5.5 identifies this group with the direct sum of the nodewise Ext-groups, so the global class determines, and is equivalently determined by, a tuple of nodewise classes $(\epsilon_{1}^{H},\dots,\epsilon_{r}^{H})$. ∎
Remark 5.2.
Proposition 5.6 should be understood as saying that the global corrected extension is assembled from localized node-to-bulk coupling data. At this level of generality we do not yet claim a basis theorem or dimension count for the individual summands; those belong to the stronger refinement developed below.
5.5. Strong nodewise Ext theorem
We now sharpen the nodewise extension-theoretic picture on the perverse side with the goal of showing that each ordinary double point contributes a one-dimensional local extension channel.
Lemma 5.7.
For each node $p_{k}\in\Sigma$, the perverse extension group
$\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\bigl(i_{k*}\mathbb{Q}_{\{p_{k}\}},IC_{X_{0}}\bigr)$
is identified with the corresponding local extension group in a sufficiently small analytic neighborhood of $p_{k}$.
Proof.
Because $i_{k*}\mathbb{Q}_{\{p_{k}\}}$ is supported at the single point $p_{k}$, every extension of $i_{k*}\mathbb{Q}_{\{p_{k}\}}$ by $IC_{X_{0}}$ is local near $p_{k}$. Restricting to a sufficiently small analytic neighborhood of $p_{k}$ therefore preserves the extension problem, and the global extension group identifies with the corresponding local perverse extension group. ∎
Proposition 5.8.
Let $p$ be an ordinary double point and let $X_{0,\mathrm{loc}}$ denote a sufficiently small analytic neighborhood of $p$ in $X_{0}$. Then the corrected local perverse ODP extension of Section 4 defines a nonzero class
$e_{\mathrm{loc}}\in\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0,\mathrm{loc}};\mathbb{Q})}\bigl(i_{*}\mathbb{Q}_{\{p\}},IC_{\mathrm{loc}}\bigr).$
Proof.
Section 4 constructs the corrected local perverse ODP extension with point-supported quotient $i_{*}\mathbb{Q}_{\{p\}}$. By construction, this extension is non-split, hence determines a nonzero class in the stated local extension group. ∎
Proposition 5.9.
There is a natural map from the local perverse extension group
$\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0,\mathrm{loc}};\mathbb{Q})}\bigl(i_{*}\mathbb{Q}_{\{p\}},IC_{\mathrm{loc}}\bigr)$
to the quotient space
$B/\operatorname{Im}(\beta),$
where
$Z(IC_{\mathrm{loc}})=(L,A,B,\alpha,\beta,\gamma)$
is the MacPherson–Vilonen zig-zag of the local intersection-complex perverse sheaf.
Proof.
Let
$0\to IC_{\mathrm{loc}}\to\mathcal{E}\to i_{*}\mathbb{Q}_{\{p\}}\to 0$
be a local perverse extension. Applying the MacPherson–Vilonen zig-zag construction yields a middle zig-zag
$Z(\mathcal{E})=(L_{E},A_{E},B_{E},\alpha_{E},\beta_{E},\gamma_{E}).$
Since the quotient $i_{*}\mathbb{Q}_{\{p\}}$ has zero open part, its zig-zag is
$Z(i_{*}\mathbb{Q}_{\{p\}})=(0,\mathbb{Q},\mathbb{Q},0,\operatorname{id},0),$
so the open part of the middle zig-zag is forced to agree with that of $IC_{\mathrm{loc}}$:
$L_{E}\cong L.$
Thus all extension data occur in the point terms $A_{E},B_{E}$ and the map $\beta_{E}$.
Choose temporary splittings of the short exact sequences of vector spaces
$0\to A\to A_{E}\to\mathbb{Q}\to 0,\qquad 0\to B\to B_{E}\to\mathbb{Q}\to 0,$
so that
$A_{E}\cong A\oplus\mathbb{Q},\qquad B_{E}\cong B\oplus\mathbb{Q}.$
With respect to these splittings, the quotient condition implies that $\beta_{E}$ induces $\operatorname{id}:\mathbb{Q}\to\mathbb{Q}$ on the quotient, hence $\beta_{E}$ has block form
$\beta_{E}=\begin{pmatrix}\beta&u\\ 0&1\end{pmatrix},\qquad u\in\mathrm{Hom}(\mathbb{Q},B)\cong B.$
We claim that the class of $u$ modulo $\operatorname{Im}(\beta)$ is independent of the chosen splittings. Indeed, replacing the chosen splitting of $A_{E}$ by one differing by $a\in\mathrm{Hom}(\mathbb{Q},A)\cong A$, and the chosen splitting of $B_{E}$ by one differing by $b\in\mathrm{Hom}(\mathbb{Q},B)\cong B$, changes the block matrix of $\beta_{E}$ by conjugation with the corresponding upper-triangular change-of-basis matrices. A direct calculation shows that the new off-diagonal term is
$u^{\prime}=u+b-\beta(a).$
In particular, changing the splitting of $A_{E}$ changes $u$ by an element of $\operatorname{Im}(\beta)$, while changing the splitting of $B_{E}$ changes the chosen representative of the same extension class in $B$. Thus the induced class
$u]\in B/\operatorname{Im}(\beta)$
does not depend on the choice of splittings.
Finally, if two perverse extensions are equivalent, then the induced isomorphism of middle terms respects the subobject $IC_{\mathrm{loc}}$ and the quotient $i_{*}\mathbb{Q}_{\{p\}}$, and therefore induces the same class $[u]$ in the quotient $B/\operatorname{Im}(\beta)$. Hence the assignment
$\mathcal{E}]\longmapsto[u]$
is well defined on extension classes. This defines the required natural map. ∎
Thus the local perverse extension class is controlled by a single quotient of the zig-zag point data, so the ordinary double point contributes at most one local extension channel.
Proposition 5.10.
One has
$\dim_{\mathbb{Q}}\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0,\mathrm{loc}};\mathbb{Q})}\bigl(i_{*}\mathbb{Q}_{\{p\}},IC_{\mathrm{loc}}\bigr)\leq 1.$
Proof.
By Proposition 5.9, a local perverse extension class determines a class in
$B/\operatorname{Im}(\beta),$
where
$Z(IC_{\mathrm{loc}})=(L,A,B,\alpha,\beta,\gamma).$
By exactness of the MacPherson–Vilonen zig-zag sequence, one has
$B/\operatorname{Im}(\beta)\cong\operatorname{Im}(\gamma)\subseteq H^{0}(i^{*}Rj_{*}j^{*}IC_{\mathrm{loc}}).$
It is therefore enough to bound the dimension of
$H^{0}(i^{*}Rj_{*}j^{*}IC_{\mathrm{loc}}).$
Now
$j^{*}IC_{\mathrm{loc}}=\mathbb{Q}_{U_{\mathrm{loc}}}[3],$
so
$H^{0}(i^{*}Rj_{*}j^{*}IC_{\mathrm{loc}})\cong H^{0}(i^{*}Rj_{*}\mathbb{Q}_{U_{\mathrm{loc}}}[3])\cong H^{3}(U_{\mathrm{loc}}\cap B;\mathbb{Q}),$
for a sufficiently small Milnor ball $B$ around the node. Here $U_{\mathrm{loc}}\cap B$ is homotopy equivalent to the local link complement, and in the ordinary double point case on a threefold this link is homotopy equivalent to $S^{2}\times S^{3}$. Consequently,
$H^{3}(U_{\mathrm{loc}}\cap B;\mathbb{Q})\cong\mathbb{Q}.$
In particular,
$\dim_{\mathbb{Q}}H^{0}(i^{*}Rj_{*}j^{*}IC_{\mathrm{loc}})=1.$
Since
$B/\operatorname{Im}(\beta)\cong\operatorname{Im}(\gamma)\subseteq H^{0}(i^{*}Rj_{*}j^{*}IC_{\mathrm{loc}}),$
it follows that
$\dim_{\mathbb{Q}}(B/\operatorname{Im}(\beta))\leq 1.$
Therefore the local perverse extension group also has dimension at most one. ∎
Corollary 5.11.
One has
$\dim_{\mathbb{Q}}\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0,\mathrm{loc}};\mathbb{Q})}\bigl(i_{*}\mathbb{Q}_{\{p\}},IC_{\mathrm{loc}}\bigr)=1,$
and the local corrected perverse ODP class $e_{\mathrm{loc}}$ is its generator.
Proof.
By Proposition 5.8, the local extension group is nonzero. By Proposition 5.10, it has dimension at most one. Hence it is one-dimensional, and the nonzero class $e_{\mathrm{loc}}$ is a generator. ∎
Corollary 5.12.
For each node $p_{k}\in\Sigma$, one has
$\dim_{\mathbb{Q}}\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\bigl(i_{k*}\mathbb{Q}_{\{p_{k}\}},IC_{X_{0}}\bigr)=1.$
Moreover, the image of the local corrected perverse ODP class defines a generator
$e_{k}\in\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\bigl(i_{k*}\mathbb{Q}_{\{p_{k}\}},IC_{X_{0}}\bigr).$
Proof.
Apply Lemma 5.7 together with Corollary 5.11. ∎
Definition 5.13.
For each node $p_{k}\in\Sigma$, let
$e_{k}\in\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\bigl(i_{k*}\mathbb{Q}_{\{p_{k}\}},IC_{X_{0}}\bigr)$
denote the generator defined by the local corrected perverse ODP extension.
Corollary 5.14.
There is a nodewise decomposition
$\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\bigl(Q_{\Sigma},IC_{X_{0}}\bigr)\cong\bigoplus_{k=1}^{r}\mathbb{Q}\,e_{k},\qquad Q_{\Sigma}:=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}}.$
In particular, the node set $\Sigma$ indexes a distinguished basis of local extension channels on the perverse side.
Proof.
This follows from Lemma 5.4 and Corollary 5.12. ∎
Proposition 5.15.
The global corrected perverse extension class is determined by a unique expansion
$\mathcal{P}]_{\mathrm{perv}}=\sum_{k=1}^{r}c_{k}\,e_{k}$
for uniquely determined coefficients $c_{k}\in\mathbb{Q}$.
Proof.
By Corollary 5.14, the global perverse extension space has basis $\{e_{1},\dots,e_{r}\}$. Hence the global corrected perverse extension class admits a unique expansion in that basis. ∎
Remark 5.3.
Corollary 5.14 and Proposition 5.15 give the first explicit mathematical shadow of the later domain-wall interpretation: each node contributes a distinguished one-dimensional local coupling channel, and the global corrected class is encoded by the coefficient vector
$(c_{1},\dots,c_{r}).$
A fuller Hodge-theoretic and limiting-mixed-Hodge-structure interpretation of these coefficients belongs to a stronger refinement of the theory.
5.6. Realization and compatibility with the perverse extension space
Applying the realization functor to the finite point-supported quotient object of Proposition 5.3 gives
$\mathrm{rat}(Q_{\Sigma}^{H})\cong\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}}.$
Thus the node-wise singular quotient in the mixed-Hodge-module setting is the precise Hodge-theoretic refinement of the finite direct sum of skyscraper perverse sheaves appearing on the perverse side.
Moreover, exactness of realization induces a natural map
$\mathrm{Ext}^{1}_{MHM(X_{0})}\bigl(Q_{\Sigma}^{H},IC^{H}_{X_{0}}\bigr)\longrightarrow\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\bigl(Q_{\Sigma},IC_{X_{0}}\bigr),$
where
$Q_{\Sigma}:=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}}.$
Under Lemma 5.4, the target decomposes as
$\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\bigl(Q_{\Sigma},IC_{X_{0}}\bigr)\cong\bigoplus_{k=1}^{r}\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\bigl(i_{k*}\mathbb{Q}_{\{p_{k}\}},IC_{X_{0}}\bigr).$
Hence the global mixed-Hodge-module extension to be constructed in the next section is a genuine refinement of the finite multi-node corrected perverse extension
$0\to IC_{X_{0}}\to\mathcal{P}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}}\to 0.$
Remark 5.4.
The decomposition above gives a mathematically precise version of the later domain-wall picture: the point-supported node sectors are localized degrees of freedom, while the extension classes in the nodewise summands measure their couplings to the bulk intersection-complex sector.
6. Global gluing theorem
We now pass from the local ordinary double point model to the global finite multi-node setting. Section 4 constructs the local mixed-Hodge-module correction block at a single node, while Section 5 identifies the finite point-supported singular quotient object
$Q_{\Sigma}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
The remaining task is to assemble the local gluing data over all nodes and thereby construct the global corrected mixed Hodge module on $X_{0}$.
6.1. The global singular object
Let
$\Sigma=\{p_{1},\dots,p_{r}\}\subset X_{0}$
be the finite set of ordinary double points of the central fiber, and let
$U:=X_{0}\setminus\Sigma,\qquad j:U\hookrightarrow X_{0},\qquad i_{k}:\{p_{k}\}\hookrightarrow X_{0}$
be the corresponding inclusions. Set
$\mathcal{M}^{\prime}_{U}:=\mathbb{Q}^{H}_{U}[3],\qquad\mathcal{M}^{\prime\prime}_{\Sigma}:=Q_{\Sigma}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Lemma 6.1.
The finite family of local singular objects
$\{i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\}_{k=1}^{r}$
defines canonically the global point-supported mixed Hodge module
$\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Proof.
This is exactly the finite point-supported mixed Hodge module of Proposition 5.2. Since the node set $\Sigma$ is a finite disjoint union of points, the corresponding point-supported mixed Hodge modules have disjoint support and therefore form a canonical finite direct sum in $MHM(X_{0})$. ∎
6.2. Assembly of the local gluing morphisms
For each node $p_{k}$, Section 4 provides a local point-supported block
$W_{k}^{H}=i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
together with local gluing morphisms
$u_{k}:\psi_{\pi,1}^{H}(\mathcal{M}^{\prime}_{U})\longrightarrow W_{k}^{H},\qquad v_{k}:W_{k}^{H}\longrightarrow\psi_{\pi,1}^{H}(\mathcal{M}^{\prime}_{U})(-1),$
satisfying
$v_{k}u_{k}=N_{k}$
on the corresponding local ordinary-double-point sector.
Proposition 6.2.
The local gluing morphisms assemble to global morphisms
$u_{\Sigma}:\psi_{\pi,1}^{H}(\mathcal{M}^{\prime}_{U})\longrightarrow\mathcal{M}^{\prime\prime}_{\Sigma},\qquad v_{\Sigma}:\mathcal{M}^{\prime\prime}_{\Sigma}\longrightarrow\psi_{\pi,1}^{H}(\mathcal{M}^{\prime}_{U})(-1).$
Proof.
For each $k$, the local map $u_{k}$ has target $W_{k}^{H}$. Since
$\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}W_{k}^{H},$
the universal property of the direct sum yields a unique morphism
$u_{\Sigma}:\psi_{\pi,1}^{H}(\mathcal{M}^{\prime}_{U})\to\mathcal{M}^{\prime\prime}_{\Sigma}$
whose $k$-th component is $u_{k}$.
Similarly, each $v_{k}$ has source $W_{k}^{H}$, so the universal property of the direct sum yields a unique morphism
$v_{\Sigma}:\mathcal{M}^{\prime\prime}_{\Sigma}\to\psi_{\pi,1}^{H}(\mathcal{M}^{\prime}_{U})(-1)$
whose restriction to the $k$-th summand is $v_{k}$. ∎
Proposition 6.3.
The global morphisms $u_{\Sigma}$ and $v_{\Sigma}$ satisfy
$v_{\Sigma}u_{\Sigma}=N$
on the direct sum of the local ordinary-double-point vanishing sectors of $\psi_{\pi,1}^{H}(\mathcal{M}^{\prime}_{U})$.
Proof.
For each node $p_{k}$, Section 4 gives the normalized relation
$v_{k}u_{k}=N_{k}$
on the corresponding local ODP sector. Because the node supports are pairwise disjoint and
$\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}W_{k}^{H},$
the assembled maps satisfy
$v_{\Sigma}u_{\Sigma}=\sum_{k=1}^{r}v_{k}u_{k}=\sum_{k=1}^{r}N_{k}.$
On the finite node-supported ordinary-double-point sector, the global nilpotent monodromy operator is exactly the sum of these local nilpotent operators, hence
$v_{\Sigma}u_{\Sigma}=N.$
6.3. The global gluing datum
The assembled datum $(M^{\prime}_{U},M^{\prime\prime}_{\Sigma},u_{\Sigma},v_{\Sigma})$ satisfies the precise compatibility required for Saito’s principal-divisor gluing theorem.
Theorem 6.4 (Global gluing datum).
The data
$(\mathcal{M}^{\prime}_{U},\mathcal{M}^{\prime\prime}_{\Sigma},u_{\Sigma},v_{\Sigma})$
form a valid divisor-gluing datum in the sense of Saito’s principal-divisor gluing theorem.
Proof.
By Lemma 6.1, the singular quotient object is $\mathcal{M}^{\prime\prime}_{\Sigma}$. By Proposition 6.2, the local gluing maps assemble to global morphisms $u_{\Sigma}$ and $v_{\Sigma}$. By Proposition 6.3, these morphisms satisfy the required relation
$v_{\Sigma}u_{\Sigma}=N$
on the finite node-supported ODP sector. Therefore the data satisfy the hypotheses of Saito’s principal-divisor gluing theorem. ∎
6.4. Construction of the global corrected mixed Hodge module
Theorem 6.5 (Global existence).
There exists an object
$\mathcal{P}^{H}\in MHM(X_{0})$
whose restriction to the smooth locus agrees with $IC^{H}_{X_{0}}|_{U}=\mathbb{Q}^{H}_{U}[3]$ and whose singular quotient is
$\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Proof.
Apply Saito’s principal-divisor gluing theorem to the global gluing datum of Theorem 6.4. This produces an object
$\mathcal{P}^{H}\in MHM(X_{0})$
with open-part restriction $\mathcal{M}^{\prime}_{U}=\mathbb{Q}^{H}_{U}[3]$ and singular quotient
$\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
The canonical middle extension of $\mathbb{Q}^{H}_{U}[3]$ across $X_{0}$ is $IC^{H}_{X_{0}}$, so $\mathcal{P}^{H}$ is the desired global corrected mixed Hodge module. ∎
6.5. Realization of the global corrected object
Theorem 6.6 (Global realization).
The realization of the global corrected mixed Hodge module is the corrected perverse object:
$\mathrm{rat}(\mathcal{P}^{H})\cong\mathcal{P}.$
Proof.
By construction, $\mathcal{P}^{H}$ is obtained by gluing the bulk object $\mathbb{Q}^{H}_{U}[3]$ to the finite point-supported singular quotient object
$\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
using the assembled gluing morphisms $u_{\Sigma},v_{\Sigma}$. Realization is exact and compatible with nearby cycles, vanishing cycles, and the gluing formalism. The realized singular quotient is
$\mathrm{rat}(\mathcal{M}^{\prime\prime}_{\Sigma})=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}},$
and the realized local gluing maps are precisely the corrected perverse local maps from Section 4. Hence the realized global gluing datum is the finite multi-node corrected perverse gluing datum, and therefore
$\mathrm{rat}(\mathcal{P}^{H})\cong\mathcal{P}.$
6.6. Global exact sequence
Lemma 6.7.
The maximal subobject of $\mathcal{P}^{H}$ whose restriction to the smooth locus is $\mathbb{Q}^{H}_{U}[3]$ is $IC^{H}_{X_{0}}$.
Proof.
By construction, $\mathcal{P}^{H}$ restricts to $\mathbb{Q}^{H}_{U}[3]$ on $U$, and its only additional singular support lies on the finite node set $\Sigma$. The canonical middle extension of $\mathbb{Q}^{H}_{U}[3]$ across $X_{0}$ with no additional singular quotient is $IC^{H}_{X_{0}}$. ∎
Theorem 6.8 (Global exact sequence).
The object $\mathcal{P}^{H}$ fits into an exact sequence
$0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\mathcal{M}^{\prime\prime}_{\Sigma}\to 0,$
that is,
$0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\to 0$
in $MHM(X_{0})$.
Proof.
By Theorem 6.5, $\mathcal{P}^{H}$ has singular quotient
$\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
By Lemma 6.7, its bulk subobject is $IC^{H}_{X_{0}}$. Therefore $\mathcal{P}^{H}$ fits into the stated exact sequence. ∎
6.7. Global cone description
Proposition 6.9 (Global cone description).
In the derived category of mixed Hodge modules on $X_{0}$, one has
$\mathcal{P}^{H}\simeq\operatorname{Cone}(v_{\Sigma})[-1].$
Proof.
The global gluing datum determines a distinguished triangle
$\mathcal{M}^{\prime\prime}_{\Sigma}\longrightarrow\psi_{\pi,1}^{H}(\mathcal{M}^{\prime}_{U})(-1)\longrightarrow\mathcal{P}^{H}[1]\overset{+1}{\longrightarrow}.$
By definition of cone in the triangulated category, the third term is
$\operatorname{Cone}(v_{\Sigma}),$
hence
$\mathcal{P}^{H}\simeq\operatorname{Cone}(v_{\Sigma})[-1].$
6.8. Global uniqueness and rigidity
Lemma 6.10.
The global extension class of $\mathcal{P}^{H}$ is determined by the collection of nodewise local extension classes.
Proof.
By Proposition 5.6, the global corrected extension class lies in
$\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{M}^{\prime\prime}_{\Sigma},IC^{H}_{X_{0}}),\qquad\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),$
and this group decomposes canonically as
$\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{M}^{\prime\prime}_{\Sigma},IC^{H}_{X_{0}})\cong\bigoplus_{k=1}^{r}\mathrm{Ext}^{1}_{MHM(X_{0})}\bigl(i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),IC^{H}_{X_{0}}\bigr).$
Thus a global extension class is equivalent to a tuple of nodewise classes, one for each node $p_{k}$. In particular, two global extensions define the same class if and only if their projections to all nodewise summands agree. ∎
Proposition 6.11 (Global rigidity).
Let $\mathcal{E}^{H}\in MHM(X_{0})$ satisfy:
  1. (1)
    $\mathrm{rat}(\mathcal{E}^{H})$ is the corrected finite multi-node perverse extension;
  2. (2)
    its singular quotient is
    $\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1);$
  3. (3)
    for each node $p_{k}\in\Sigma$, its restriction to a sufficiently small analytic neighborhood of $p_{k}$ is isomorphic to the local mixed-Hodge-module ODP extension of Theorem 4.11 (equivalently, the unique local object of Proposition 4.14).
Then $\mathcal{E}^{H}\cong\mathcal{P}^{H}$.
Proof.
Both $\mathcal{E}^{H}$ and $\mathcal{P}^{H}$ determine extension classes in
$\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{M}^{\prime\prime}_{\Sigma},IC^{H}_{X_{0}}).$
By Lemma 6.10, each such global class is completely determined by its nodewise components in the direct-sum decomposition
$\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{M}^{\prime\prime}_{\Sigma},IC^{H}_{X_{0}})\cong\bigoplus_{k=1}^{r}\mathrm{Ext}^{1}_{MHM(X_{0})}\bigl(i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),IC^{H}_{X_{0}}\bigr).$
Fix a node $p_{k}\in\Sigma$. By assumption, the restriction of $\mathcal{E}^{H}$ to a sufficiently small analytic neighborhood of $p_{k}$ is isomorphic to the local mixed-Hodge-module ODP extension of Theorem 4.11. On the other hand, the global object $\mathcal{P}^{H}$ was constructed by assembling exactly these local ODP blocks via the global gluing procedure of Section 6, so its restriction near $p_{k}$ is the same local object. By the local rigidity statement of Proposition 4.14, the corresponding local extension class is uniquely determined. Therefore the $k$-th nodewise component of the global extension class of $\mathcal{E}^{H}$ agrees with the $k$-th nodewise component of the global extension class of $\mathcal{P}^{H}$.
Since this holds for every node $p_{k}\in\Sigma$, all nodewise components of the two global classes agree. Lemma 6.10 now implies that the two global extension classes in
$\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{M}^{\prime\prime}_{\Sigma},IC^{H}_{X_{0}})$
are equal. Hence the nodewise components of the extension classes of $\mathcal{E}^{H}$ and $\mathcal{P}^{H}$ agree at every node, so by Lemma 6.10 the global extension classes agree, and therefore $\mathcal{E}^{H}\cong\mathcal{P}^{H}$. ∎
7. Hypercohomology and limiting mixed Hodge structures
In this section we extract the Hodge-theoretic consequences of the global mixed Hodge module extension constructed in Section 6. The key point is that the same nearby-cycle mixed-Hodge-module formalism controls both the corrected object $\mathcal{P}^{H}$ and the limiting mixed Hodge structure of the degeneration. The singular quotient
$\mathcal{V}^{H}:=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
is therefore not merely a point-supported correction term on the central fiber: it is the precise mixed-Hodge-module refinement of the local vanishing contribution that appears in the limiting mixed Hodge structure.
7.1. Hypercohomology of nearby-cycle mixed Hodge modules
Let
$F=\mathbb{Q}_{\mathcal{X}}[3].$
By Saito’s theory, the nearby-cycle object
$\psi_{\pi}^{H}(F)$
is a mixed Hodge module on $X_{0}$, and its hypercohomology carries the limiting mixed Hodge structure associated with the degeneration. More precisely, applying hypercohomology to the nearby-cycle mixed-Hodge-module formalism yields the limiting mixed Hodge structure on the cohomology of the nearby fiber in the sense of Schmid and Steenbrink [16, 18].
Lemma 7.1.
For each $m$, the mixed Hodge structure
$\mathbb{H}^{m}(X_{0},\psi_{\pi}^{H}(F))$
is the limiting mixed Hodge structure associated with the degeneration in degree $m$.
Proof.
This is the standard compatibility of nearby cycles in Saito’s category with the classical limiting mixed Hodge structure formalism. ∎
7.2. The point-supported quotient and the vanishing sector
By Theorem 6.8, the global corrected mixed Hodge module fits into an exact sequence
$0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\mathcal{V}^{H}\to 0,\qquad\mathcal{V}^{H}:=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Thus $\mathcal{V}^{H}$ is exactly the singular quotient object identified in Sections 5 and 6.
Proposition 7.2.
The quotient
$\mathcal{V}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
realizes the rank-$r$ local vanishing contribution in the nearby-cycle formalism.
Proof.
At each node $p_{k}$, the local Milnor fiber has the homotopy type of $S^{3}$, so the reduced cohomology is rank one in degree $3$ and vanishes otherwise. By the local construction of Section 4, the corresponding mixed-Hodge-module refinement of the local singular quotient is
$i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Summing over the finite node set gives the global quotient $\mathcal{V}^{H}$. Since nearby and vanishing cycles in $MHM$ refine the corresponding perverse objects and their realization is compatible with the local variation picture, this quotient is exactly the mixed-Hodge-module realization of the rank-$r$ local vanishing sector. ∎
Corollary 7.3.
Applying the realization functor to $\mathcal{V}^{H}$ yields
$\mathrm{rat}(\mathcal{V}^{H})\cong\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}},$
the singular quotient in the corrected perverse extension.
Proof.
This follows immediately from the exactness of $\mathrm{rat}$ and the identity
$\mathrm{rat}(i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1))\cong i_{k*}\mathbb{Q}_{\{p_{k}\}}.$
7.3. Hypercohomology / LMHS comparison theorem
Write again
$\mathcal{V}^{H}:=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
By Theorem 6.8, the corrected mixed-Hodge-module object fits into an exact sequence
(7.1) $0\longrightarrow IC^{H}_{X_{0}}\longrightarrow\mathcal{P}^{H}\longrightarrow\mathcal{V}^{H}\longrightarrow 0.$
Applying hypercohomology yields a long exact sequence in mixed Hodge structures
(7.2) $\cdots\to\mathbb{H}^{m}(X_{0},IC^{H}_{X_{0}})\longrightarrow\mathbb{H}^{m}(X_{0},\mathcal{P}^{H})\longrightarrow\mathbb{H}^{m}(X_{0},\mathcal{V}^{H})\overset{\partial_{m}}{\longrightarrow}\mathbb{H}^{m+1}(X_{0},IC^{H}_{X_{0}})\to\cdots.$
The mixed-Hodge-module extension in Eqn. (7.1) first produces an extension on hypercohomology in the category of mixed Hodge structures, and the claim is that this induced extension is exactly the finite vanishing-sector extension coming from the LMHS nearby-cycle formalism. By Proposition 6.9,
$\mathcal{P}^{H}\simeq\operatorname{Cone}(v_{\Sigma})[-1].$
Thus $\mathcal{P}^{H}$ is the global object produced from the same nearby-cycle data that carry the LMHS.
Definition 7.4.
The LMHS extension class of the finite multi-node degeneration is the extension class in the category of mixed Hodge structures obtained by applying hypercohomology to the global corrected mixed-Hodge-module extension Eqn. (7.1) and restricting to the rank-$r$ local vanishing sector carried by $\mathcal{V}^{H}$.
Lemma 7.5.
There is a commutative diagram of long exact sequences in mixed Hodge structures
$\mathbb{H}^{m}(X_{0},IC^{H}_{X_{0}})$$\mathbb{H}^{m}(X_{0},\mathcal{P}^{H})$$\mathbb{H}^{m}(X_{0},\mathcal{V}^{H})$$\mathbb{H}^{m+1}(X_{0},IC^{H}_{X_{0}})$$\mathbb{H}^{m}(X_{0},IC^{H}_{X_{0}})$$\mathbb{H}^{m}(X_{0},\psi_{\pi}^{H}(F))$$\mathbb{H}^{m}(X_{0},\mathcal{V}^{H})$$\mathbb{H}^{m+1}(X_{0},IC^{H}_{X_{0}})$$\scriptstyle{\partial_{m}}$
whose bottom row is the nearby-cycle long exact sequence restricted to the finite vanishing sector.
Proof.
The top row is obtained from Eqn. (7.1). The bottom row is obtained from the nearby-cycle formalism after restricting to the finite point-supported vanishing contribution $\mathcal{V}^{H}$. Since $\mathcal{P}^{H}$ is the global corrected object produced from the same nearby-cycle and variation data, and
$\mathcal{P}^{H}\simeq\operatorname{Cone}(v_{\Sigma})[-1]$
by Proposition 6.9, the two constructions come from the same distinguished triangle in the derived category of mixed Hodge modules. Therefore the associated long exact sequences agree, yielding the commutative diagram. ∎
Theorem 7.6.
The extension induced by Eqn. (7.1) on hypercohomology is canonically the LMHS extension class of Definition 7.4. Equivalently, the point-supported quotient
$\mathcal{V}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
is the precise mixed-Hodge-module source of the rank-$r$ local vanishing extension in the limiting mixed Hodge structure.
Proof.
By Lemma 7.5, the connecting morphisms in the long exact sequence (7.2) coincide with the connecting morphisms in the nearby-cycle long exact sequence on the finite vanishing sector. By Definition 7.4, the latter determine the LMHS extension class. Therefore the extension induced by (7.1) on hypercohomology is canonically the LMHS extension class. ∎
7.4. Extension classes and comparison questions
Let
$\mathcal{P}^{H}]\in\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{V}^{H},IC^{H}_{X_{0}})$
denote the Yoneda class of Eqn. (7.1). Applying the exact realization functor to Eqn. (7.1) yields the perverse exact sequence
(7.3) $0\longrightarrow IC_{X_{0}}\longrightarrow\mathcal{P}\longrightarrow\mathcal{V}\longrightarrow 0,\qquad\mathcal{V}:=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}},$
with Yoneda class
$\mathcal{P}]\in\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}(\mathcal{V},IC_{X_{0}}).$
Lemma 7.7.
The hypercohomology functor sends the Yoneda class $\mathcal{P}^{H}]$ to the extension class in $MHS$ determined by the connecting morphisms of the long exact sequence (7.2).
Proof.
This is the standard compatibility of a cohomological functor with Yoneda extensions in the heart of a $t$-structure: applying hypercohomology to Eqn. (7.1) yields the long exact sequence Eqn. (7.2), and the connecting morphisms encode the induced extension class in mixed Hodge structures. ∎
Proposition 7.8.
There are natural maps
$\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{V}^{H},IC^{H}_{X_{0}})\overset{\mathrm{rat}}{\longrightarrow}\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}(\mathcal{V},IC_{X_{0}})$
and
$\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{V}^{H},IC^{H}_{X_{0}})\overset{\mathbb{H}^{\bullet}}{\longrightarrow}\mathrm{Ext}^{1}_{MHS}\!\bigl(\mathbb{H}^{\bullet}(X_{0},\mathcal{V}^{H}),\mathbb{H}^{\bullet}(X_{0},IC^{H}_{X_{0}})\bigr),$
and the class $\mathcal{P}^{H}]$ maps to the perverse extension class $\mathcal{P}]$ under the first map and to the LMHS extension class under the second.
Proof.
The first map is induced by exactness of $\mathrm{rat}$. The second is induced by hypercohomology together with Lemma 7.7. By Theorem 7.6, the induced hypercohomological extension is canonically the LMHS extension class. ∎
Theorem 7.9.
The corrected extension is organized simultaneously at three levels:
  1. (1)
    as the mixed-Hodge-module extension class
    $\mathcal{P}^{H}]\in\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{V}^{H},IC^{H}_{X_{0}});$
  2. (2)
    as its realized perverse extension class
    $\mathcal{P}]\in\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}(\mathcal{V},IC_{X_{0}});$
  3. (3)
    as its induced LMHS extension class on hypercohomology in $MHS$.
Moreover, realization and hypercohomology relate these classes functorially, and all three arise from the same nearby-cycle and variation data of the degeneration.
Proof.
The relation between the first two classes is given by exactness of $\mathrm{rat}$. The relation between the first and third is given by Lemma 7.7 and Theorem 7.6. The final assertion follows from the construction of $\mathcal{P}^{H}$ from the nearby-cycle and variation data in Sections 4–6. ∎
7.5. Weight filtrations and local vanishing pieces
We conclude by recording the weight-theoretic normalization relevant to the point-supported quotient. By the conventions fixed earlier, the local vanishing contribution attached to a node is represented by
$i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
The Tate twist (1)(-1) is the standard integral Tate twist in the category of mixed Hodge modules and mixed Hodge structures, and it is compatible with the action of the nilpotent monodromy operator
$N:\psi_{\pi,1}^{H}(\mathcal{M})\to\psi_{\pi,1}^{H}(\mathcal{M})(-1).$
In particular, the local singular quotient is normalized so that it agrees with the monodromy-theoretic shift built into Saito’s nearby-cycle formalism. The present paper does not attempt a full analysis of the weight filtration on all of $\mathcal{P}^{H}$ or on all hypercohomology groups
$\mathbb{H}^{m}(X_{0},\mathcal{P}^{H}).$
What is established is the precise weight-normalized form of the point-supported quotient and its role as the Hodge-theoretic refinement of the local vanishing sector. A fuller analysis of the weight filtration on the global object and its relation to stronger Kähler-package phenomena lies beyond the scope of the present paper.
8. Auxiliary structural results
This section gathers several structural consequences and auxiliary viewpoints that clarify the meaning of the mixed-Hodge-module extension constructed in the previous sections. None of the results below is needed for the existence or realization theorems themselves, but each of them helps explain how the global object $\mathcal{P}^{H}$ fits simultaneously into the perverse-sheaf, mixed-Hodge-module, and later physical pictures.
8.1. Recollement and perverse-side extension groups
Let
$X_{0}=U\sqcup\Sigma,\qquad\Sigma=\{p_{1},\dots,p_{r}\},$
with
$j:U\hookrightarrow X_{0},\qquad i_{k}:\{p_{k}\}\hookrightarrow X_{0}.$
The category $\mathrm{Perv}(X_{0};\mathbb{Q})$ admits the standard recollement description associated with this open–closed decomposition [2, 7, 4]. In particular, every perverse extension of $\mathbb{Q}_{U}[3]$ is controlled by its restriction to the smooth locus together with a point-supported contribution on $\Sigma$.
In the single-node case, [11] shows that the corrected perverse object fits into a canonical short exact sequence
$0\to IC_{X_{0}}\to\mathcal{P}\to i_{*}\mathbb{Q}_{\{p\}}\to 0,$
and that the corresponding extension class lies in
$\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}(i_{*}\mathbb{Q}_{\{p\}},IC_{X_{0}}).$
The finite multi-node case is the natural direct-sum generalization of that picture. The global perverse extension
$0\to IC_{X_{0}}\to\mathcal{P}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}}\to 0$
is classified by an element of
$\mathrm{Ext}^{1}_{\mathrm{Perv}(X_{0};\mathbb{Q})}\Bigl(\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}},IC_{X_{0}}\Bigr).$
Section 5 strengthens this description by isolating a distinguished nodewise family of perverse extension classes and, in the ordinary double point case, a basis
$\{e_{1},\dots,e_{r}\}$
for the corresponding nodewise extension space. Thus the global corrected perverse extension class admits a concrete nodewise expansion
$\mathcal{P}]_{\mathrm{perv}}=\sum_{k=1}^{r}c_{k}\,e_{k},$
which records how the individual node sectors are assembled into a single corrected perverse object.
The mixed-Hodge-module extension constructed in the present paper is a refinement of exactly this perverse extension class. More precisely, applying the realization functor to the exact sequence
$0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\to 0$
recovers the perverse-side extension class in the recollement category. Thus the present theorem package should be viewed as an internal Hodge-theoretic lift of the same extension-theoretic structure already visible in $\mathrm{Perv}(X_{0};\mathbb{Q})$.
8.2. Compatibility with Banagl–Budur–Maxim
A useful methodological precedent for the present construction is the work of Banagl–Budur–Maxim [1]. In their setting, one starts with a perverse sheaf constructed from nearby-cycle data for an isolated hypersurface singularity and shows that it underlies a mixed Hodge module, so that its hypercohomology carries canonical mixed Hodge structures. Their object is not the corrected perverse extension considered here, but the formal pattern is closely related: a perverse object arising from degeneration data is shown to admit a Hodge-theoretic refinement internal to Saito’s theory.
The present paper differs from [1] in both geometric focus and theorem content. The object we refine is the canonical corrected perverse extension $\mathcal{P}$ attached to a conifold degeneration with finitely many ordinary double points, rather than the intersection-space complex attached to an isolated hypersurface singularity. Nonetheless, the conceptual similarity is important. In both settings:
  • nearby and vanishing cycles provide the underlying singular contribution;
  • the resulting perverse object is not arbitrary but canonically attached to the degeneration;
  • the mixed-Hodge-module formalism is the correct category in which to internalize the Hodge-theoretic structure.
Accordingly, [1] should be viewed not as a proof source for the specific multi-node conifold theorem established here, but as a methodological model for the kind of refinement the present paper carries out explicitly in the ordinary double point setting.
8.3. Quiver-theoretic shadow of the multi-node extension
The finite multi-node exact sequence
$0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\to 0$
suggests a natural algebraic shadow. Each node contributes a rank-one point-supported summand, and the global extension class records how these local summands are coupled to the bulk object $IC^{H}_{X_{0}}$. Section 5 makes this more explicit on the perverse side: the nodewise extension space carries a distinguished basis
$\{e_{1},\dots,e_{r}\},$
and the corrected global perverse class expands as
$\mathcal{P}]_{\mathrm{perv}}=\sum_{k=1}^{r}c_{k}\,e_{k}.$
This coefficient vector already provides the first algebraic shadow of a finite interaction graph or quiver attached to the degeneration. The present paper does not attempt a full quiver-theoretic classification of the mixed-Hodge-module extension, but the theorem package established above makes such a picture mathematically natural.
The data
$\bigl(IC^{H}_{X_{0}},\{i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\}_{k=1}^{r},[\mathcal{P}^{H}]\bigr)$
thus has the formal structure of a bulk object together with finitely many localized rank-one objects and a global extension class describing their coupling. This is the natural precursor of any later quiver, schober, or wall-crossing refinement.
8.4. Toward the domain-wall interpretation
The theorem package of the present paper gives a precise mathematical foundation for the bulk/localized-sector language proposed in the later physical interpretation. The object
$IC^{H}_{X_{0}}$
plays the role of the bulk geometric sector: it is the Hodge-module refinement of the intersection complex of the singular fiber and therefore carries the part of the degeneration that persists away from the nodes. By contrast, the quotient
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
is supported entirely at the singular points and records the rank-one local vanishing contribution of each collapsing cycle.
The exact sequence
$0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\to 0$
therefore admits the following rigorous interpretation: it is a bulk/localized-sector coupling law in the category of mixed Hodge modules. The nontriviality of the extension class means that the localized vanishing-cycle sectors cannot be split off from the bulk geometry without changing the global mixed-Hodge-module structure of the degeneration. After applying realization, one recovers the corresponding perverse extension
$0\to IC_{X_{0}}\to\mathcal{P}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}_{\{p_{k}\}}\to 0,$
which is exactly the sheaf-theoretic structure that later physical work interprets as the coupling of localized framed sectors to a bulk geometric sector.
The significance of the present paper is that this interpretation no longer rests only on the perverse side. Theorems 6.8 and 7.6 show that the same localized quotient also contributes the vanishing part of the limiting mixed Hodge structure on hypercohomology. Thus the bulk/localized-sector interpretation is simultaneously visible in the mixed-Hodge-module extension, in its perverse realization, and in the Hodge-theoretic degeneration data. This is the precise mathematical content that later physical applications may safely use.
9. Consequences and further directions
9.1. What has been proved
The main result of the present paper is the construction of a fully internal mixed-Hodge-module refinement of the canonical corrected perverse object for a finite multi-node conifold degeneration. More precisely, for a one-parameter degeneration whose central fiber $X_{0}$ has ordinary double points
$\Sigma=\{p_{1},\dots,p_{r}\},$
we constructed an object
$\mathcal{P}^{H}\in MHM(X_{0})$
whose realization is the corrected perverse object
$\mathcal{P}.$
This gives a theorem-level Hodge-theoretic refinement of the canonical corrected perverse extension in the finite multi-node ordinary double point setting.
The construction proceeds by combining four essential ingredients. First, the local ordinary double point model yields a rank-one point-supported mixed-Hodge-module block at each singular point. Second, the finite support decomposition identifies the global singular contribution as the direct sum of these local rank-one blocks. Third, Saito’s divisor-case gluing formalism provides the mechanism for assembling the smooth-locus Hodge-module data and the finite node contributions into a single global object on $X_{0}$. Fourth, realization carries the resulting mixed-Hodge-module extension to the canonical corrected perverse object, while hypercohomology identifies the same point-supported quotient with the vanishing sector of the limiting mixed Hodge structure.
The resulting theorem package may be summarized as follows:
  • there exists an object
    $\mathcal{P}^{H}\in MHM(X_{0});$
  • its realization satisfies
    $\mathrm{rat}(\mathcal{P}^{H})\cong\mathcal{P};$
  • it fits into an exact sequence
    $0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\to 0$
    in $MHM(X_{0})$;
  • the point-supported quotient realizes the rank-$r$ local vanishing contribution of the degeneration;
  • the hypercohomology of $\mathcal{P}^{H}$ is functorially related to the limiting mixed Hodge structure through the same nearby-cycle mixed-Hodge-module formalism;
  • and the corrected extension is organized simultaneously as a class in mixed Hodge modules, as its realized class in perverse sheaves, and as an induced extension class on hypercohomology in mixed Hodge structures.
In particular, the present paper upgrades the corrected perverse extension from a constructible or perverse object to a genuine mixed-Hodge-module object carrying internal Hodge-theoretic structure. This is precisely the refinement that was identified as missing in the earlier nearby-cycle bridge picture and that was not available in the original perverse-sheaf construction.
9.2. Consequences for subsequent work
The significance of the present paper is not only that it proves a new mixed-Hodge-module theorem, but that it changes the status of the larger program built around the corrected perverse object. Before the present construction, one could compare the corrected perverse extension and the Hodge-theoretic degeneration data only at the level of a common nearby-cycle origin. After the present paper, one has an actual object
$\mathcal{P}^{H}\in MHM(X_{0})$
whose realization is the corrected perverse extension and whose point-supported quotient is the Hodge-theoretic refinement of the local vanishing sector.
This has several immediate consequences.
First, it supplies a theorem-level foundation for the study of global extension data in finite multi-node degenerations. The corrected extension is no longer merely an object of $\mathrm{Perv}(X_{0};\mathbb{Q})$; it is now an extension in the category of mixed Hodge modules, with an internal Hodge-theoretic quotient and an induced extension on hypercohomology. Thus later work may safely treat the global extension data as genuinely Hodge-theoretic rather than merely perverse.
Second, the extension space itself is now substantially better understood. The singular quotient is the finite direct sum
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),$
and the corresponding global extension class decomposes nodewise. On the perverse side, the finite ordinary double point setting yields distinguished nodewise extension classes
$e_{1},\dots,e_{r},$
so that the corrected perverse extension admits a concrete nodewise organization. This is the first precise mathematical form of the statement that the degeneration contains finitely many localized sectors coupled to a bulk geometric sector.
Third, it gives a rigorous mathematical substrate for the quiver-theoretic, schober-theoretic, and wall-crossing directions anticipated earlier. The finite collection of rank-one point-supported objects
$\{i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\}_{k=1}^{r}$
together with the global extension class of $\mathcal{P}^{H}$ already has the formal shape of a finite system of localized sectors coupled to a bulk geometric sector. This is exactly the kind of structure from which quiver, schober, and wall-crossing formalisms are expected to emerge.
Fourth, it provides theorem-level mathematical support for subsequent physical interpretations of multi-node conifold degenerations. In particular, the exact sequence
$0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\to 0$
gives a precise mixed-Hodge-module version of the bulk/localized-sector decomposition that later work may interpret in physical language. The important point is that this interpretation no longer rests on heuristic analogy alone: the localized quotient is now an actual theorem-level object in $MHM(X_{0})$, and its contribution to the limiting mixed Hodge structure is part of the proved hypercohomological picture.
In this sense, the present paper should be viewed as the technical keystone between the earlier perverse-sheaf construction and the later categorical and physical developments. It does not yet solve the full wall-crossing or schober problem, but it supplies the foundational Hodge-theoretic object that those later theories require.
9.3. Further directions
The theorem package established here resolves the finite multi-node ordinary double point case, but it also points directly to a number of natural next problems.
(1) Stronger comparison of extension classes.
The present paper now relates the mixed-Hodge-module extension, the realized perverse extension, and the induced extension on hypercohomology. A natural next step is to strengthen this further into a more explicit comparison theorem among extension classes in $MHM(X_{0})$, in $\mathrm{Perv}(X_{0};\mathbb{Q})$, and in the category of mixed Hodge structures, with sharper control of the maps between these extension spaces and of the corresponding nodewise generators.
(2) Hodge-theoretic lifts of the nodewise Ext generators.
Section 5 isolates distinguished nodewise extension classes on the perverse side. A natural next problem is to construct and understand corresponding classes directly in
$\mathrm{Ext}^{1}_{MHM(X_{0})}\!\left(i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),IC^{H}_{X_{0}}\right)$
and to compare them functorially with the perverse and hypercohomological extension classes. Such a refinement would sharpen the rigidity theory and make the Hodge-theoretic content of the global extension more explicit.
(3) Arbitrary stratified singular loci.
The most immediate geometric extension is to pass beyond finite collections of ordinary double points and treat degenerations whose singular locus has higher-dimensional strata. At the perverse level, the corrected object in such a setting is expected to differ from the intersection complex by a constructible singular contribution supported on the singular locus. The mixed-Hodge-module analogue would require replacing the finite direct sum of point-supported rank-one objects by a more general singular quotient, together with a corresponding extension of the divisor-gluing construction carried out here.
(4) Kähler-package questions for $\mathbb{H}^{*}(X_{0},\mathcal{P}^{H})$.
Once the mixed-Hodge-module refinement $\mathcal{P}^{H}$ is available, it becomes meaningful to ask whether its hypercohomology satisfies analogues of the classical Kähler-package properties known for intersection cohomology: duality, Lefschetz-type isomorphisms, and Hodge–Riemann bilinear relations. The present paper establishes the Hodge-theoretic object to which such questions may now be addressed, but does not attempt to resolve them.
(5) Quiver, schober, and wall-crossing refinements.
The finite-node extension data already suggests a quiver-theoretic organization of the localized node contributions and their coupling to the bulk sector. Likewise, the relation of the corrected perverse object to spherical-monodromy phenomena suggests that a categorified schober-theoretic version of the present construction should exist. The mixed-Hodge-module refinement proved here is the natural precursor of those higher-categorical structures.
(6) Full LMHS and coefficient interpretation.
The present paper identifies the point-supported quotient as the source of the finite vanishing sector in the limiting mixed Hodge structure. A natural next question is to understand more fully the coefficients with which the global corrected class is assembled from the nodewise sectors and to give those coefficients a direct Hodge-theoretic and limiting-mixed-Hodge-theoretic interpretation.
Taken together, these directions show that the finite multi-node ordinary double point theorem proved here is not the end of the story, but the first fully internal Hodge-theoretic foundation for a broader program linking nearby cycles, mixed Hodge modules, canonical perverse extensions, extension spaces, quiver-type structures, and physical wall-crossing phenomena.